-- SimonaPaiano - 14 Oct 2020

INAF queue observing, 2020 Oct 13 - Oct 14

Observer: A. Cardwell
Telescope Operator: Josh
Astronomers from Italy: F. Cusano, S. Paiano, A. Rossi
Instrument: LUCI-BINO

Summary

Due to the covid-19 pandemic situation these observations have been performed by LBTO personel
through the indications of LBT Italian observers (F.Cusano, S. Paiano) connected from Italy.

At the sunset the sky is clear.

Night duration (12 degree el twilight) 1:50 - 12:30 UT = 10 h 40m

Observing time: 10 h 19 min (96.7%)

Weather time loss: * 0 h 0 min ( 0 %)*

technical time loss * 0h 21 min (3.3 %)*

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Images NotesSorted ascending
Cassara' dark LUCI 1-2        
Paiano dark LUCI 1-2        
Feruglio dark LUCI 1 -2        
Tell_CassarĂ  HIP13917 LUCI 1,2 zJ200   luci1.0119-0122, luci2.0030-003  
Schmidt J0823+3427 LBC blue+red r-band, i-band 0.55+0.55 hrs   1 OB done
Polletta flat LUCI 1 z-band     for PLCK_G67.1+22.5
CassarĂ  flat,arcs LUCI 1,2 zJ200     for VIPE1
Rossi_BACKUP GRB161007A LBC blue+red g,r on LBCblue + z LBCred 0.67+0.67 hrs   LBC imaging completed
Calibration Skyflats LBCblue+red r+z     note: there are good flats in both PA maybe just for z-band
Polletta PLCK_G67.1+22.5 LUCI 1 z - ESM 1.5 hrs   seeing = 0.6" - 0.85" - 1.3" - 0.85" good images from 64
Feruglio S82X _1940 - A6 LUCI 1 G210-H - ESM 0.7 hrs luci1.0103-113 seeing=0.65", PA angle not correct and possible tilt. Probably the target was not in the slit.
D'Onofrio NGC1042 LBC blue+red g-band, r-band 1.6+1.6 hrs hr seeing ~1.", completed
Tell_Feruglio HIP 115428 LUCI 1 G210-H - ESM     Telluric - A6
CassarĂ  VIPE1 LUCI 1,2 zJ200 0.8+0.8hrs luci1.0127-0140, luci2.0038.0051 We cannot see the science spectra or emission lines. Problem with the mask ?
NOTES:

Polletta run

The exposures 0062-0063 must to be rejected (see details in the Andrew's log).

Some elepticity in luci1.0073

Variable seeing, up to 1.3", at 03:15-03:25UT. At 03:38 it dropped to 0.8"

Feruglio run

After carefully measurting the angle on PanStarrs images and verifying the target (luckily visible after background subtraction) we do the hypothesis that the PA angle of S82X _1940 - A6 may be no correct and the target is out of the slit.

Cassara' run

to-update

Details

detailed log by A. Cardwell at this link: https://docs.google.com/document/d/1sis_Y_biPPjm7VkoRA9gsvRfmn7sW_U-jIMI_QEYbXQ/edit?usp=sharing
Topic revision: r1 - 31 Jul 2024, UnknownUser
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback